Hi, please help me solve this problem, thank you.
The question is:
Indicate an equation that satisfies the description of each parabole. y=x2
a) A parabole with a summit of (0, -3)
b) A parabole with the summit of (5,1)
c) A parabole with a summit of (4,0)
d) A parabole with a summit of (3,5)
e)A parabole with a summit of (-8,6)

Answer by stanbon(75887)

Indicate an equation that satisfies the description of each parabole. y=x^2
a) A parabola with a summit of (0, -3)
y=-x^2 -3
---------------
b) A parabola with the summit of (5,1)
y-1=-(x-5)^2 or y=-(x-5)^2+1
---------------
c) A parabola with a summit of (4,0)
y=-(x-4)^2
------------------
d) A parabola with a summit of (3,5)
y-5=-(x-3)^2 or y=-(x-3)^2+5
-----------------
e)A parabola with a summit of (-8,6)
y-6=-(x+8)^2 or y=-(x+8)^2+6
------------
Cheers,
Stan H.